OpenStudy (ty-quan19):

What is the discriminant of the quadratic equation 2x + 5x2 = 1?

jhonyy9 (jhonyy9):

2x +5x^2 = 1 so the quadratic equation will be 5x^2 +2x -1 = 0 formula for discriminant is D = b^2 -4ac so in this case D = 2^2 -4(5)*(-1) = 4 -20*(-1) = 4+20 = 24 so the discriminant is equal 24 hope helped and is understandably